OpenStudy (anonymous):

Can anyone verify 1-tan^2(x/2) = (2cosx)/(1+cosx)?

OpenStudy (mertsj):

\[\tan ^2\frac{x}{2}=\frac{1-\cos x}{1+\cos x}\]

OpenStudy (mertsj):

Might try that substitution ans see where it leads.

OpenStudy (anonymous):

I did that and now im stuck.

OpenStudy (mertsj):

\[1=\frac{1+\cos x}{1+\cos x}\]

OpenStudy (mertsj):

Make that substitution as well.

OpenStudy (anonymous):

Where?

OpenStudy (mertsj):

On the left side.

OpenStudy (anonymous):

but i thought i had 1-cosx/1+cosx

OpenStudy (mertsj):